Consider a monopolistically competitive firm facing an indirect demand curve given by P = 20 – 2Q. If the profit-maximizing price is 16, what is the marginal cost?

a) $14

b) $10

c) $12

d) cannot be determined from the information given.
Answer #1

The correct option is C.) $12

At the profit-maximizing level for a monopolistically competitive firm,

MR = MC

P = 20 -2Q

At P = 16, Q= 2

TR = P.Q

     = (20-2Q).(Q)

    = 20Q - 2Q^2

MR= dTR/dQ = 20 -4Q

MR = 20 - 4(2)

    = 12

Thus, MC = 12